determine weather the graphs of the equations are perpendicular
4x-3y=8
4y-3x=7

Answer by subudear(62)

4x-3y=8 .....(1)
4y-3x=7 .....(2)
slope for first line y = (4/3)x - 8/3 is m1 = 4/3
slope for second line y = (3/4)x + 7/4 is m2 = 3/4
for the lines to be perpendicular m1*m2 should be equal to -1
m1*m2 = (4/3)*(3/4) = 1
so these lines are not perpendicular